OpenStudy (anonymous):

need help i will fan and medal !! Find the area of the parallelogram. 14 square units 48 square units 24 square units 28 square units

OpenStudy (cbarredo1):

Is there an image or something?

OpenStudy (wolf1728):

We need a graphic

OpenStudy (cbarredo1):

48 square units

OpenStudy (anonymous):

can you explain how ?

OpenStudy (cbarredo1):

A=b x h the base would be six and the height 8, just count the squares :)

OpenStudy (cbarredo1):

Count from D to C and then count the y axis were the two lines cross

OpenStudy (anonymous):

oh okay ! thank you !!!

OpenStudy (wolf1728):

Area of a parallelogram = base * height base = 6 height^2 = 8^2 - 3^2 height^2 = 64 -9 height^2 = 55 height = 7.4162 Area = base * height Area = 44.497

OpenStudy (wolf1728):

CBARREDO1 got it right my answer is wrong :-(

OpenStudy (anonymous):

i appreciate the help still !@wolf1728

OpenStudy (cbarredo1):

;)

OpenStudy (wolf1728):

okay - I'm glad I didn't mislead you. :-)